romano cycles from her house to school at 15 miles per hour. upon arriving, she realizes that it is Saturday and immediately cycles home at 23 miles per hour. if the entire round trip takes her 32 minutes, what is her average speed, in miles per hour, for the entire round trip?
Answer by mananth(16946)

Let the distance to school be x miles
speed to school = 15 mph
time taken to go to school = x/15
..
Similarly time to return from school = x/23
Total time taken = 32 minutes.= 32/60 hours
x/15 + x/23 = 32/60
23x+15x / 15*23 = 32/60
38x = 32*15*23 /60
38x=184
x= 4.84 miles
UP & down she travels 2*4.84 miles = 9.68
Total time taken = 32/60 hours
Average speed = 9.68 / 32/60
9.68 * 60 /32
18.15 mph